OpenStudy (anonymous):

A large bag of sugar weighs 7 one over two lb. How much does one and one fourth bags of sugar weigh? :7/1/8 8/3/4 9/3/4 9/3/8

OpenStudy (texaschic101):

1/7.5 = 1.25/x (1 bag to 7 1/2 lbs = 1 1/4 bag to x lbs) cross multiply x = (1.25)(7.5) x = 9.375 = 9 3/8 lbs

OpenStudy (anonymous):

yAY I Was rIGHT Thanks @texaschic101 I gave you a medal